OpenStudy (anonymous):

Graph each pair of parametric equations. x = 3 sin^3t y = 3 cos^3t

OpenStudy (anonymous):

@agent0smith ?

OpenStudy (anonymous):

@primeralph

OpenStudy (dumbcow):

try to put "y" in terms of "x" to graph on a x,y axis solve x(t) for "t" \[t = \sin^{-1} (\sqrt[3]{\frac{x}{3}})\] plug into y(t) \[y = 3\cos ^{3} [\sin^{-1} (\sqrt[3]{\frac{x}{3}})] = 3(\sqrt{1-(\frac{x}{3})^{2/3}})^{3}\]

OpenStudy (anonymous):

im very confused

OpenStudy (anonymous):

i dont even understnd what you did from the start

OpenStudy (anonymous):

how did you get the inverse sin..and all f that

OpenStudy (anonymous):

@dumbcow ...?